OpenStudy (anonymous):

How do you write an equation in slope-intercept form for this problem?! (0,0); y=-3x+2

OpenStudy (the_fizicx99):

This is already in slope intercept from >.<

OpenStudy (anonymous):

yea as fizi said, y = mx + b :) cheers

OpenStudy (anonymous):

already in the form

OpenStudy (anonymous):

@tHe_FiZiCx99 I realized that too but in my book it says "write an equation in slope-intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation." So I think it means I need to substitute the x and y values but their both 0

OpenStudy (the_fizicx99):

Ohh ^^ well what're the "given points" and its perpendicular so its the reciprocal, so suppose 3/4x I'd be -4/3x

OpenStudy (anonymous):

So would it be y=-1/3x+2 ?

OpenStudy (the_fizicx99):

what're the given points it says it gives you. You would plug in y - y1=m(x-x1)

OpenStudy (anonymous):

The given points are (0,0)

OpenStudy (anonymous):

Did I write the answer right? y=-1/3x

OpenStudy (the_fizicx99):

For example suppose it says, find the equation that is perpendicular to y = 3/4x + 6 and the given points are (-8,6) So the perp if 3/4 is -4/3x So, to find the equation with the new given points plug them into, y - y1=m(x-x1) So, y-6=-4/3(x-(-8)) -4/3 * 8/1 --> -4 * 8 = -32/3 Y-6 = -4/3x - 10.6 +6 +6 y = -4/3x -4.6

OpenStudy (the_fizicx99):

of*

OpenStudy (anonymous):

Thanks! I got it now

OpenStudy (the_fizicx99):

Yw bo ;*